# WISE: Web Information Satire and Fakeness Evaluation

### v1

@@ Abstract.
Distinguishing fake or untrue news from satire or humor poses a unique challenge due to their overlapping linguistic features and divergent intent. This study develops WISE (Web Information Satire and Fakeness Evaluation) framework which benchmarks eight lightweight transformer models alongside two baseline models on a balanced dataset of 20,000 samples from Fakeddit, annotated as either fake news or satire. Using stratified 5-fold cross-validation, we evaluate models across comprehensive metrics including accuracy, precision, recall, F1-score, ROC-AUC, PR-AUC, MCC, Brier score, and Expected Calibration Error. Our evaluation reveals that MiniLM, a lightweight model, achieves the highest accuracy (87.58%) among all models, while RoBERTa-base achieves the highest ROC-AUC (95.42%) and strong accuracy (87.36%). DistilBERT offers an excellent efficiency-accuracy trade-off with 86.28% accuracy and 93.90% ROC-AUC. Statistical tests confirm significant performance differences between models, with paired t-tests and McNemar tests providing rigorous comparisons. Our findings highlight that lightweight models can match or exceed baseline performance, offering actionable insights for deploying misinformation detection systems in real-world, resource-constrained settings.
@@
copyright: acmlicensedjournalyear: 2025doi: TBDconference: The 19th ACM International Conference on Web Search and Data Mining ; February 22–26, 2026; Boise, IDisbn: 978-1-4503-XXXX-X/20XX/XXccs: Computing methodologies Natural language processingccs: Theory of computation Machine learning theoryccs: Information systems Information retrieval
@@ 1. Introduction
The rapid proliferation of online information has intensified the need to distinguish malicious fake news from satirical content. Fake news, defined as intentionally false information presented as factual, has become a major societal concern, often conflated with other misleading content such as satire and factual errors (golbeck2018fake). False news spreads faster and farther than truthful news, largely driven by human behavior rather than automation (vosoughi2018spread; aral2018wired), and can cause serious harm, as seen in conspiracy-driven vandalism incidents. Addressing misinformation requires avoiding the conflation of fake news with satire, as the latter mimics news style while differing in intent. Classifiers trained to separate satire from real news often also detect fake news, reflecting overlapping linguistic patterns (horne2017fake_satire_similarity), yet satire aims for humor and critique rather than deception. Mislabeling satire undermines its role in public discourse, where irony and exaggeration are used for social commentary. Thus, automated systems must reliably separate harmful misinformation from benign satire to protect both information integrity and satirical expression.
@@
Fake news detection has evolved from feature-based classifiers using linguistic and stylistic cues (e.g., n-grams, readability, sentiment) to deep learning methods. Early approaches with support vector machines (SVMs) or Naive Bayes on handcrafted features achieved moderate success (rastogi2023fake_news_survey). Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) improved representation learning, and transformer-based models like Bidirectional Encoder Representations from Transformers (BERT) further advanced accuracy via contextualized understanding (kaliyar2021fakebert). However, most research frames detection as real vs. fake or as part of broader multi-class misinformation taxonomies, overlooking the fake-vs-satire challenge. Satirical writing employs rhetorical devices such as irony and sarcasm, complicating detection. Only a small body of work, such as Levi et al. (levi2019fake_vs_satire), addresses this distinction using semantic, linguistic, and coherence features with contextual models. Multimodal satire detection incorporating visual cues (e.g., ViLBERT) shows potential (nguyen2021multimodal), but there remains limited evaluation of lightweight transformers on fake-vs-satire tasks. This gap limits understanding of whether compact models can capture these nuances while remaining efficient—an essential consideration for real-world deployments.
@@
In this study, we benchmark eight lightweight transformer models alongside two baseline models for distinguishing fake news from satire. We use a balanced subset of the Fakeddit dataset (nakamura2019fakeddit), comprising 20,000 samples (10,000 satire and 10,000 fake news), ensuring robust evaluation through stratified 5-fold cross-validation. We assess comprehensive classification metrics including accuracy, precision, recall, F1-score (macro and per-class), ROC-AUC, PR-AUC, Matthews Correlation Coefficient (MCC), Brier score, and Expected Calibration Error (ECE), alongside training and inference efficiency. Our contributions are: (1) the first systematic benchmark of multiple lightweight transformers on a large-scale fake-vs-satire dataset using rigorous cross-validation, (2) comprehensive evaluation with full statistical testing including paired t-tests, McNemar tests, and DeLong tests, (3) analysis of how depth, parameter sharing, and distillation influence detection of satirical vs. deceptive cues, and (4) deployment-oriented insights for building high-accuracy, resource-efficient misinformation detection systems.
@@ 2.1. Fake News Detection
Fake news detection techniques have progressed markedly over the past decade. Early approaches relied on handcrafted linguistic features, for example, n-gram frequency patterns, lexical complexity, and sentiment cues, to distinguish deceptive content (potthast2018stylometric). Such methods showed that deceptive writing often exhibits telltale stylistic signs, such as negative affect or unusual phrasing, that can be captured via bag-of-words or parse-tree representations (rubin2016fake). As the field advanced, machine learning and deep neural networks gained prominence. Researchers began applying convolutional and recurrent neural networks to automatically learn features from text, yielding better accuracy than earlier feature-engineered models (zhang2020overview). For instance, long short-term memory and convolutional architectures, sometimes combined with metadata, outperformed traditional classifiers on benchmark fake news datasets (ruchansky2017csi). The emergence of transformer-based language models like BERT has since revolutionized NLP, and several studies report that fine-tuning such models significantly boosts fake news classification performance (kaliyar2021fakebert; jawahar2020bertstructure). BERT-based detectors, such as FakeBERT, can capture subtle linguistic nuances and contextual cues, leading to state-of-the-art results in identifying misinformation (kaliyar2021fakebert). However, most of this literature frames fake news detection as a binary classification, distinguishing only between real and fake news (rubin2015deception). Relatively little work has addressed finer-grained distinctions between different kinds of misleading content. Notably, Rubin (rubin2015deception) identified categories like outright fabrications, large-scale hoaxes, and satire as separate challenges, highlighting that current systems typically do not attempt to differentiate satire from intentional fake news. This gap motivates specialized investigations into sub-types of misinformation.
@@ 2.2. Satire Detection
In contrast to general fake news detection, the automatic recognition of satirical news has received comparatively less scholarly attention. Satire is a form of deceptive news that uses humor, irony, and exaggeration not to mislead maliciously, but to entertain or critique (bourgonje2017clickbait; mihalcea2006humor). These traits make satire especially challenging for algorithms to identify, as the language often mimics legitimate news while embedding subtle cues such as absurd scenarios or sarcastic tone that signal its true intent (burfoot2009satire). Early studies approached satire detection through linguistic and stylistic features. For example, Burfoot and Baldwin (burfoot2009satire) and later Rubin et al. (rubin2016satire_detection) used features like absurdity, slang, negative affect, and punctuation patterns combined with support vector machine classifiers to successfully distinguish satirical articles. Rubin’s system achieved an F-score around 0.87 by leveraging such handcrafted cues (rubin2016satire_detection). Some subsequent works experimented with deep learning. Yang et al. (yang2017humor) proposed a hierarchical attention network to capture satire at both paragraph and article levels. However, the application of advanced neural architectures to satire-specific detection remains limited. Only a few datasets exist for this task; one prominent example is the Golbeck et al. (golbeck2018fake) dataset of online news labeled as fake or satire. While this resource has enabled initial studies on satirical news classification, the use of transformer-based architectures such as BERT or its variants for satire detection is still in its early stages. Research has only begun to explore whether these models can reliably capture the nuanced humor and irony in satire, and current findings suggest it remains an open problem requiring further investigation.
@@ 2.3. Lightweight Transformer Models
The powerful performance of large transformers such as BERT comes at the cost of high model complexity and substantial computational resource requirements. This challenge has motivated the development of lightweight transformer models that preserve strong language understanding capabilities while operating more efficiently. One prominent approach is knowledge distillation, as demonstrated by TinyBERT, which compresses BERT into a smaller model by transferring knowledge from a large teacher model to a compact student model (jiao2020tinybert). A four-layer TinyBERT can achieve approximately 96–97% of its teacher’s accuracy on General Language Understanding Evaluation (GLUE) benchmarks while being more than seven times smaller and nearly nine times faster during inference (jiao2020tinybert). Another strategy is exemplified by ALBERT (A Lite BERT), which employs parameter-reduction techniques such as factorized embedding layers and cross-layer parameter sharing to drastically reduce the number of parameters without significantly impacting performance (lan2020albert). ALBERT achieved new state-of-the-art results on benchmarks like Stanford Question Answering Dataset (SQuAD) and GLUE with 18 times fewer parameters than BERT-large. Similarly, MiniLM uses a deep self-attention distillation technique focusing on key transformer attention components, enabling the student model to retain over 99% of BERT’s performance on tasks such as question answering and natural language inference while using only about half the model size and computation (Wang et al., 2020). Additional compact models include DistilBERT, which is approximately 40 percemt smaller and 60% faster than BERT-base while maintaining around 95–97% of its language understanding capability (sanh2019distilbert), and MobileBERT, which optimizes BERT for mobile devices through a bottleneck architecture (sun2020mobilebert). These lightweight transformers have demonstrated strong performance across diverse NLP tasks, showing that substantial model compression is possible with minimal accuracy loss, making them attractive for deployment in resource-constrained or real-time fake news detection systems.
@@ 2.4. Benchmark Studies
There have been numerous benchmark studies evaluating transformer models on general NLP tasks, yet none have specifically addressed the challenge of distinguishing fake news from satire. General benchmark frameworks such as GLUE and SuperGLUE have facilitated systematic comparisons of models including BERT, RoBERTa, and ALBERT on tasks such as text classification, entailment, and question answering (wang2019glue; wang2020minilm). These evaluations have driven progress in the field by identifying the architectures that perform best under various conditions. However, the satire-versus-fake news classification problem has largely been absent from these benchmark suites. A notable exception is the dataset introduced by Golbeck et al. (golbeck2018fake), which contains 283 fake news articles and 203 satire articles within the same political domain, each with verified labels. This dataset demonstrated strong inter-annotator agreement (approx.= 0.69) and 84% labeling accuracy, making it a reliable resource for satire detection research. It has been used in several studies to test classification models, with reported baseline accuracies generally in the mid-70% to low-80% range for distinguishing satire from fake news (das2020heuristic). Another dataset, FakeNewsNet and its derivative Fakeddit, includes satire as one of multiple misinformation categories (shu2020fakenewsnet; nakov2021clef). However, these datasets are not exclusively focused on satire-vs-fake classification and often include social media context that differs from mainstream news. Overall, there is a clear lack of comprehensive benchmarks dedicated to this specific classification task. The present study addresses this gap by systematically evaluating multiple lightweight transformer models on the satire-versus-fake news distinction, providing the first in-depth comparison of their performance in this specialized misinformation detection domain.
@@ 3.1. Dataset
We used a balanced subset of the Fakeddit dataset (nakamura2019fakeddit), a large-scale multimodal dataset for fake news detection collected from Reddit. Fakeddit contains posts labeled across six categories: True, Satire/Parody, Misleading Content, Manipulated Content, False Connection, and Imposter Content. For our binary classification task, we curated a balanced dataset focusing on the distinction between satire and fake news.
@@ 3.3.2. Preprocessing and Tokenization
Preprocessing preserves linguistic integrity while preparing inputs for each model’s tokenizer. Text samples are normalized with consistent whitespace and special character handling, then tokenized with each model’s native tokenizer from the Hugging Face Transformers library. Inputs are capped at 256 tokens (configurable via max_length), which comfortably covers the dataset’s text length distribution (mean 82.7 characters, range 51-296). This token limit ensures efficient processing while maintaining sufficient context for classification. Tokenized inputs are cached to disk to avoid redundant tokenization across folds and models, improving experimental efficiency.
@@ 3.3.3. Training Configuration
All experiments run with PyTorch on macOS using MPS (Metal Performance Shaders) acceleration under a shared hardware setup. Hyperparameters are fixed across all models for fair comparison, as detailed in Table 3. We use the AdamW optimizer with ReduceLROnPlateau scheduler, reducing the learning rate by half when validation loss plateaus. Early stopping is based on validation loss to prevent overfitting, with the best model checkpoint saved based on validation performance. The primary metric for model selection is macro-averaged F1-score, though early stopping uses validation loss for better generalization. All models are trained with the same random seed (42) for reproducibility, with deterministic operations enabled where possible.

### v2

@@ Abstract.
Distinguishing fake or untrue news from satire or humor poses a unique challenge due to their overlapping linguistic features and divergent intent. This study develops WISE (Web Information Satire and Fakeness Evaluation) framework, which benchmarks eight lightweight transformer models alongside two baseline models on a balanced dataset of 20,000 samples from Fakeddit, annotated as either fake news or satire. Using stratified 5-fold cross-validation, we evaluate models across comprehensive metrics including accuracy, precision, recall, F1-score, ROC-AUC, PR-AUC, MCC, Brier score, and Expected Calibration Error. Our evaluation reveals that MiniLM, a lightweight model, achieves the highest accuracy (87.58%) among all models, while RoBERTa-base achieves the highest ROC-AUC (95.42%) and strong accuracy (87.36%). DistilBERT offers an excellent efficiency-accuracy trade-off with 86.28% accuracy and 93.90% ROC-AUC. Statistical tests confirm significant performance differences between models, with paired t-tests and McNemar tests providing rigorous comparisons. Our findings highlight that lightweight models can match or exceed baseline performance, offering actionable insights for deploying misinformation detection systems in real-world, resource-constrained settings.
@@
journalyear: 2026copyright: ccconference: The Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: The Nineteenth ACM International Conference on Web Search and Data Mining (WSDM Companion ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3779211.3793174isbn: 979-8-4007-2358-2/2026/02ccs: Computing methodologies Natural language processingccs: Theory of computation Machine learning theoryccs: Information systems Information retrieval
Figure 1. Overview of the WISE framework
@@ 1. Introduction
The rapid proliferation of online information has intensified the need to distinguish malicious fake news from satirical content. Fake news, defined as intentionally false information presented as factual, has become a major societal concern, often conflated with other misleading content such as satire and factual errors (Golbeck et al., 2018). False news spreads faster and farther than truthful news, largely driven by human behavior rather than automation (Vosoughi et al., 2018; Aral et al., 2018), and can cause serious harm, as seen in conspiracy-driven vandalism incidents. Addressing misinformation requires avoiding the conflation of fake news with satire, as the latter mimics news style while differing in intent. Classifiers trained to separate satire from real news often also detect fake news, reflecting overlapping linguistic patterns (Horne and Adali, 2017), yet satire aims for humor and critique rather than deception. Mislabeling satire or emotion associated with it undermine its role in public discourse, where irony and exaggeration are used for social commentary (Das and Zubaidi, 2023; Das, 2023). Thus, automated systems must reliably separate harmful misinformation from benign satire to protect both information integrity and satirical expression.
@@
Fake news detection has evolved from feature-based classifiers using linguistic and stylistic cues (e.g., n-grams, readability, sentiment) to deep learning methods. Early approaches with support vector machines (SVMs) or Naive Bayes on handcrafted features achieved moderate success (Rastogi, 2023). Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) improved representation learning, and transformer-based models like Bidirectional Encoder Representations from Transformers (BERT) further advanced accuracy via contextualized understanding (Kaliyar et al., 2021). However, most research frames detection as real vs. fake or as part of broader multi-class misinformation taxonomies, overlooking the fake-vs-satire challenge. Satirical writing employs rhetorical devices such as irony and sarcasm, complicating detection. Only a small body of work, such as Levi et al. (Levi et al., 2019), addresses this distinction using semantic, linguistic, and coherence features with contextual models. Multimodal satire detection incorporating visual cues shows potential (Nguyen and Pham, 2021), but there remains limited evaluation of lightweight transformers on fake-vs-satire tasks. This gap limits understanding of whether compact models can capture these nuances while remaining efficient—an essential consideration for real-world deployments.
@@
In this study, we benchmark eight lightweight transformer models alongside two baseline models for distinguishing fake news from satire. We use a balanced subset of the Fakeddit dataset (Nakamura et al., 2019), comprising 20,000 samples (10,000 satire and 10,000 fake news), ensuring robust evaluation through stratified 5-fold cross-validation. We assess comprehensive classification metrics including accuracy, precision, recall, F1-score (macro and per-class), Receiver Operating Characteristic Area Under the Curve (ROC-AUC), Precision-Recall Area Under the Curve (PR-AUC), Matthews Correlation Coefficient (MCC), Brier score, and Expected Calibration Error (ECE), alongside training and inference efficiency. Our contributions are: (1) the first systematic benchmark of multiple lightweight transformers on a large-scale fake-vs-satire dataset using rigorous cross-validation, (2) comprehensive evaluation with full statistical testing including paired t-tests, McNemar tests, and DeLong tests, (3) analysis of how depth, parameter sharing, and distillation influence detection of satirical vs. deceptive cues, and (4) deployment-oriented insights for building high-accuracy, resource-efficient misinformation detection systems.
@@ 2.1. Fake News Detection
Fake news detection techniques have progressed markedly over the past decade. Early approaches relied on handcrafted linguistic features, for example, n-gram frequency patterns, lexical complexity, and sentiment cues, to distinguish deceptive content (Potthast et al., 2018). Such methods showed that deceptive writing often exhibits telltale stylistic signs, such as negative affect or unusual phrasing, that can be captured via bag-of-words or parse-tree representations (Rubin et al., 2016). As the field advanced, machine learning and deep neural networks gained prominence. Researchers began applying convolutional and recurrent neural networks to automatically learn features from text, yielding better accuracy than earlier feature-engineered models (Zhang and Ghorbani, 2020). For instance, long short-term memory and convolutional architectures, sometimes combined with metadata, outperformed traditional classifiers on benchmark fake news datasets (Ruchansky et al., 2017). The emergence of transformer-based language models like BERT has since revolutionized NLP, and several studies report that fine-tuning such models significantly boosts fake news classification performance (Kaliyar et al., 2021; Jawahar et al., 2020). BERT-based detectors, such as FakeBERT, can capture subtle linguistic nuances and contextual cues, leading to state-of-the-art results in identifying misinformation (Kaliyar et al., 2021). However, most of this literature frames fake news detection as a binary classification, distinguishing only between real and fake news (Rubin, 2015). Relatively little work has addressed finer-grained distinctions between different kinds of misleading content. Notably, Rubin (Rubin, 2015) identified categories like outright fabrications, large-scale hoaxes, and satire as separate challenges, highlighting that current systems typically do not attempt to differentiate satire from intentional fake news. This gap motivates specialized investigations into sub-types of misinformation.
@@ 2.2. Satire Detection
In contrast to general fake news detection, the automatic recognition of satirical news has received comparatively less scholarly attention. Satire is a form of deceptive news that uses humor, irony, and exaggeration not to mislead maliciously, but to entertain or critique (Bourgonje et al., 2017; Mihalcea and Strapparava, 2006). These traits make satire especially challenging for algorithms to identify, as the language often mimics legitimate news while embedding subtle cues such as absurd scenarios or sarcastic tone that signal its true intent (Burfoot and Baldwin, 2009). Early studies approached satire detection through linguistic and stylistic features. For example, Burfoot and Baldwin (Burfoot and Baldwin, 2009) and later Rubin et al. (Rubin and others, 2016) used features like absurdity, slang, negative affect, and punctuation patterns combined with support vector machine classifiers to successfully distinguish satirical articles. Rubin’s system achieved an F-score around 0.87 by leveraging such handcrafted cues (Rubin and others, 2016). Some subsequent works experimented with deep learning. Yang et al. (Yang et al., 2017) proposed a hierarchical attention network to capture satire at both paragraph and article levels. However, the application of advanced neural architectures to satire-specific detection remains limited. Only a few datasets exist for this task; one prominent example is the Golbeck et al. (Golbeck et al., 2018) dataset of online news labeled as fake or satire. While this resource has enabled initial studies on satirical news classification, the use of transformer-based architectures such as BERT or its variants for satire detection is still in its early stages. Research has only begun to explore whether these models can reliably capture the nuanced humor and irony in satire, and current findings suggest it remains an open problem requiring further investigation (Seth and Sharaff, 2024).
@@ 2.3. Lightweight Transformer Models
The powerful performance of large transformers such as BERT comes at the cost of high model complexity and substantial computational resource requirements. This challenge has motivated the development of lightweight transformer models that preserve strong language understanding capabilities while operating more efficiently. One prominent approach is knowledge distillation, as demonstrated by TinyBERT, which compresses BERT into a smaller model by transferring knowledge from a large teacher model to a compact student model (Jiao et al., 2020). A four-layer TinyBERT can achieve approximately 96–97% of its teacher’s accuracy on General Language Understanding Evaluation (GLUE) benchmarks while being more than seven times smaller and nearly nine times faster during inference (Jiao et al., 2020). Another strategy is exemplified by ALBERT (A Lite BERT), which employs parameter-reduction techniques such as factorized embedding layers and cross-layer parameter sharing to drastically reduce the number of parameters without significantly impacting performance (Lan et al., 2020). ALBERT achieved new state-of-the-art results on benchmarks like Stanford Question Answering Dataset (SQuAD) and GLUE with 18 times fewer parameters than BERT-large. Similarly, MiniLM uses a deep self-attention distillation technique focusing on key transformer attention components, enabling the student model to retain over 99% of BERT’s performance on tasks such as question answering and natural language inference while using only about half the model size and computation (Wang et al., 2019). Additional compact models include DistilBERT, which is approximately 40 percemt smaller and 60% faster than BERT-base while maintaining around 95–97% of its language understanding capability (Sanh et al., 2019), and MobileBERT, which optimizes BERT for mobile devices through a bottleneck architecture (Sun et al., 2020). These lightweight transformers have demonstrated strong performance across diverse NLP tasks, showing that substantial model compression is possible with minimal accuracy loss, making them attractive for deployment in resource-constrained or real-time fake news detection systems.
@@ 2.4. Benchmark Studies
There have been numerous benchmark studies evaluating transformer models on general NLP tasks, yet none have specifically addressed the challenge of distinguishing fake news from satire. General benchmark frameworks such as GLUE and SuperGLUE have facilitated systematic comparisons of models including BERT, RoBERTa, and ALBERT on tasks such as text classification, entailment, and question answering (Wang et al., 2019, 2020). These evaluations have driven progress in the field by identifying the architectures that perform best under various conditions. However, the satire-versus-fake news classification problem has largely been absent from these benchmark suites. A notable exception is the dataset introduced by Golbeck et al. (Golbeck et al., 2018), which contains 283 fake news articles and 203 satire articles within the same political domain, each with verified labels. This dataset demonstrated strong inter-annotator agreement (approx.= 0.69) and 84% labeling accuracy, making it a reliable resource for satire detection research. It has been used in several studies to test classification models, with reported baseline accuracies generally in the mid-70% to low-80% range for distinguishing satire from fake news (Das et al., 2020). Another dataset, FakeNewsNet and its derivative Fakeddit, includes satire as one of multiple misinformation categories (Shu et al., 2020; Nakov et al., 2021). However, these datasets are not exclusively focused on satire-vs-fake classification and often include social media context that differs from mainstream news (Das, 2024). Overall, there is a clear lack of comprehensive benchmarks dedicated to this specific classification task. The present study addresses this gap by systematically evaluating multiple lightweight transformer models on the satire-versus-fake news distinction, providing the first in-depth comparison of their performance in this specialized misinformation detection domain.
@@ 3.1. Dataset
We used a balanced subset of the Fakeddit dataset (Nakamura et al., 2019), a large-scale multimodal dataset for fake news detection collected from Reddit. Fakeddit contains posts labeled across six categories: True, Satire/Parody, Misleading Content, Manipulated Content, False Connection, and Imposter Content. For our binary classification task, we curated a balanced dataset focusing on the distinction between satire and fake news.
@@ 3.3.2. Preprocessing and Tokenization
Preprocessing preserves linguistic integrity while preparing inputs for each model’s tokenizer. Text samples are normalized with consistent whitespace and special character handling, then tokenized with each model’s native tokenizer from the Hugging Face Transformers library (Wolf et al., 2020). Inputs are capped at 256 tokens (configurable via max_length), which comfortably covers the dataset’s text length distribution (mean 82.7 characters, range 51-296). This token limit ensures efficient processing while maintaining sufficient context for classification. Tokenized inputs are cached to disk to avoid redundant tokenization across folds and models, improving experimental efficiency.
@@ 3.3.3. Training Configuration
All experiments run with PyTorch (Paszke et al., 2019) on macOS using MPS (Metal Performance Shaders) acceleration under a shared hardware setup. Hyperparameters are fixed across all models for fair comparison, as detailed in Table 3. We use the AdamW (Loshchilov and Hutter, 2017) optimizer with ReduceLROnPlateau scheduler, reducing the learning rate by half when validation loss plateaus. Early stopping is based on validation loss to prevent overfitting, with the best model checkpoint saved based on validation performance. The primary metric for model selection is macro-averaged F1-score, though early stopping uses validation loss for better generalization. All models are trained with the same random seed (42) for reproducibility, with deterministic operations enabled where possible.